central angles and intercepted arcs

with the central angle 30 degrees, what is the measure of the intercepted arc in terms of pie in a circle with a radius of 10 cm
Answer #1
If circle of radius 30 cm intercepts an arc of 6 cm is assumed then


A central angle is measured by its intercepted arc.
Let's denote the length of the intercepted arc with s, and the length of the radius r. So,
s = 6 cm and r = 30 cm.
When a central angle intercepts an arc whose length measure equals the length measure of the radius of the circle, this central angle has a measure 1radian.
To find the angle in our problem we use the following relationship:
measure of an angle in radians = (length of the intercepted arc)/(length of the radius)
measure of our angle = s/r = 6/30 = 1/5 radians.
Now, we need to convert this measure angle in radians to degrees.
Since pi radians = 180 degrees, then
1 radians = 180/pi degrees, so:
1/5 radians = (1/5)(180/pi) degrees = 36/pi degrees, or approximate to 11.5 degrees.

answered by: Shainah
Answer #2
A central angle is measured by its intercepted arc.
Let's denote the length of the intercepted arc with s, and the length of the radius r. So,
s = 6 cm and r = 30 cm.
When a central angle intercepts an arc whose length measure equals the length measure of the radius of the circle, this central angle has a measure 1radian.
To find the angle in our problem we use the following relationship:
measure of an angle in radians = (length of the intercepted arc)/(length of the radius)
measure of our angle = s/r = 6/30 = 1/5 radians.
Now, we need to convert this measure angle in radians to degrees.
Since pi radians = 180 degrees, then
1 radians = 180/pi degrees, so:
1/5 radians = (1/5)(180/pi) degrees = 36/pi degrees, or approximate to 11.5 degrees.
answered by: Destine'e
Answer #3
l= 30/360*2pir= 5.24 cm
answered by: Xcequiel
Answer #4

10*π/6= 5π/3cms

answered by: yann alexandria
Answer #5
A central angle is measured by its intercepted arc.
Let's denote the length of the intercepted arc with s, and the length of the radius r. So,
s = 6 cm and r = 30 cm.
When a central angle intercepts an arc whose length measure equals the length measure of the radius of the circle, this central angle has a measure 1radian.
To find the angle in our problem we use the following relationship:
measure of an angle in radians = (length of the intercepted arc)/(length of the radius)
measure of our angle = s/r = 6/30 = 1/5 radians.
Now, we need to convert this measure angle in radians to degrees.
Since pi radians = 180 degrees, then
1 radians = 180/pi degrees, so:
1/5 radians = (1/5)(180/pi) degrees = 36/pi degrees, or approximate to 11.5 degrees.

Read more:http://wiki.answers.com/Q/A_central_angle_of_a_circle_of_radius_30_cm_intercepts_an_arc_of_6_cm_Express_the_central_angle_in_radians_and_in_degrees#ixzz1lIMVX96z